# Permutation twisted cohomology, remixed

### v2

@@ 5. The (re-)twisted cohomology ring
Let $\mathbb{N}^{\mathcal{B}(G)}=\{q\colon\mathcal{B}(G)\to\mathbb{N}\}$ be the monoid of twists, i.e. tuples of non-negative integers indexed by the $\mathcal{B}(G)$. Equivalently, the monoid of twists identifies with the submonoid of $\operatorname{Pic}(\mathscr{K}(G))$ generated by $\mathcal{B}(G)$. Consider the $(\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)})$-graded ring
@@
Its multiplication is induced by the tensor product in $\mathscr{K}(G)$. Note that only non-positive shifts $s\leq 0$ produce non-zero homomorphisms. We call $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ the twisted cohomology ring of $G$. It is convenient to write
@@
This construction may not immediately adapt if $G$ is not a $p$-group, since in this setting it is not known if one has a canonical $\mathbb{Z}$-basis for $\operatorname{Pic}(\mathscr{K}(G))$. A possible replacement could be the subgroup of $\mathscr{K}(G)$ of endotrivials arising from representation spheres for $G$. We propose that this subgroup has finite index in $\mathscr{K}(G),$ hence the analogous construction of an open cover should also hold.
To examine the $k$-vector space $\operatorname{H}^{s,q}(G)$, it suffices to choose an indecomposable representative due to the stabilizers grow property.
@@
Let $G$ be a finite $p$-group, $k(q)=C$ an effective endotrivial, and $s$ an integer. Let $f\colon k\to C[s]$ be a nonzero chain complex homomorphism. We have that $f\in\operatorname{Hom}_{\mathscr{K}(G)}(k,C[s])=0$ if and only if $\operatorname{im}(f)\subset C_{s}$ is contained in a contractible direct summand of $C$. In particular, if $C$ is indecomposable, then
@@
The converse implication is straightforward. Suppose $f\in\operatorname{Hom}_{\mathscr{K}(G)}(k,C[s])=0$, then there exists a homotopy $h\colon k\to C_{s+1}$ such that $f=d_{s+1}\circ h$. Choose a canonical $\mathbb{Z}$-basis of $C$, and consider the projection of the image of $f$ onto each indecomposable summand. Necessarily on each summand $k[G/H]$ whose projection is nonzero, the image must be the unique minimal nonzero submodule $k\subseteq k[G/H]$. Since $f=d_{s+1}\circ h$, there exists an indecomposable direct summand $M$ of $C_{s+1}$ such that $p_{k[G/H]}\circ f=d_{s+1}\circ p_{M}\circ h$. Since the stabilizers grow condition holds for $C$, the only possible homomorphisms $M\to k[G/H]$ with $k\subset M$ not in the kernel are isomorphisms, hence $M\cong k[G/H]$ and we may split off the contractible summand. An induction argument demonstrates $\operatorname{im}(f)$ is contained in a contractible direct summand of $C$ as desired. ∎
Just like in [Bal10a, BG25], we have a canonical comparison map.
@@
Same as [BG25, Proposition 13.4]. ∎
Recall the conservative functor of Theorem 1.8, the collection of modular fixed points functors $\{\hat{\Psi}^{H}\}_{H\leq G}$. For shorthand, we denote this functor by $\prod\hat{\Psi}:\mathscr{K}(G)\to\prod_{H\leq G}\operatorname{D}_{b}({G}/\!\!/{H})$. We’ll describe the “twisted” analogue of the induced homomorphism on the respective cohomology rings.
Definition 5.5.
Let $C\in\mathcal{B}(G)$, $s\in\mathbb{Z}$, and $f\colon k\to C[s]$ be a homogeneous element of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$. For every subgroup $H\leq G$, $\hat{\Psi}^{H}(C)\in\operatorname{Hom}_{\operatorname{D}_{b}({G}/\!\!/{H})}(k,\Psi^{H}(C)[s])$, and since we have an isomorphism $\Psi^{H}(C)[s]\cong k[h_{C}(H)+s]$ (determined by $\iota^{H}_{C}$), this determines an element in group cohomology $f_{H}\in\operatorname{H}^{h_{C}(H)+s}({G}/\!\!/{H},k)=\operatorname{Ext}^{h_{C}(H)+s}_{k[{G}/\!\!/{H}]}(k,k)$ determined by the fraction
$k\xrightarrow{\Psi^{H}(f)}C[s]\xleftarrow{\Psi^{H}(\iota^{H}_{C})}k[h_{C}(H)+s].$
We consider $\hat{\Psi}^{H}(f)$ an element of $\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H})$, i.e. as a morphism in $\operatorname{D}_{b}(k[{G}/\!\!/{H}])$. Note that for the endotrivial $u_{N}$ defined in [BG25, Definition 12.3] and $H=1$, the fraction $\hat{\Psi}^{1}(\iota^{G}_{u_{N}})$ is precisely the fraction $\zeta_{N}^{+}$ of [BG25, Remark 12.7]. If $H=G$, the fraction $\hat{\Psi}^{G}(\iota^{1}_{u_{N}})$ is precisely the fraction $\zeta_{N}^{-}$.
Performing this construction over all subgroups $H\leq G$ up to conjugacy determines an element $\left(\prod\hat{\Psi}\right)(f)$ in the $\operatorname{CF}(G)$-graded ring
$\operatorname{dH}^{{\scriptscriptstyle\bullet}}(G):=\prod_{H\in s_{p}(G)/G}\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H},k).$
For $\prod\hat{\Psi}$ to be a well-defined ring homomorphism, the denominator associated to an arbitrary effective endotrivial $C$ is defined as the product of $\iota^{H}_{C}$s corresponding to the unique tensor product factorization of $C$ into a product of irreducible endotrivials. This product of $\iota^{H}$s remains a quasi-isomorphism. Note that this product is well-defined up to reordering, since $h_{C}$ returns only even values for $p$ odd.
The assignment described describes the ring homomorphism $\prod\hat{\Psi}\colon\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{dH}^{\scriptscriptstyle\bullet}(G)$ induced by the conservative functor of [BG25, Theorem 7.2]. It is clear that this homomorphism is linear with respect to addition. Linearity over multiplication follows from the following roof; let $C$ and $D$ be endotrivials, $H$ a subgroup of $G$, $i$ and $j$ integers, and $f\colon k\to C[i]$ and $g\colon k\to D[j]$ homogeneous elements of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$. For shorthand, we write $\iota^{H}_{C}$ and $\iota^{H}_{D}$ for the products of $\iota^{H}_{C^{\prime}}$ and $\iota^{H}_{D^{\prime}}$ running over all irreducible endotrivial $C^{\prime}$ and $D^{\prime}$ occurring in the tensor decompositions of $C$ and $D$ respectively. Then, the following roof commutes.
[…]
Note that $\prod\hat{\Psi}$ sends tensor products of homomorphisms to Yoneda products of cohomology classes, sums of homomorphisms to Baer sums, and sends a homogeneous element in homological degree $(s,k(q))$ to a homogeneous element in degree $\underline{s}+h_{k(q)}$, where $\underline{s}$ denotes the constant superclass function returning only $s$.
Finally, we of course have a homeomorphism
$\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))\cong\bigsqcup_{H\in s_{p}(G)/G}V_{{G}/\!\!/{H}}.$
We denote the induced map on spectrum by $\prod\hat{\psi}\colon\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$, similar to the categorical case.
Proposition 5.6.
The kernel of $\prod\hat{\Psi}$ is $\otimes$-nilpotent, i.e. elements $f\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ for which $f^{\otimes n}=0$ for some $n\geq 1$.
Proof.
This follows directly from [BG25, Theorems 7.1, 7.2], since $\prod\hat{\Psi}$ is induced from the conservative family of functors $\{\hat{\Psi}^{H}\}$. ∎

### v3

@@ 5. The remixed twisted cohomology ring
Let $\mathbb{N}^{\mathcal{B}(G)}=\{q\colon\mathcal{B}(G)\to\mathbb{N}\}$ be the monoid of twists, i.e., tuples of non-negative integers indexed by elements of $\mathcal{B}(G)$. Equivalently, the monoid of twists identifies with the submonoid of $\operatorname{Pic}(\mathscr{K}(G))$ generated by $\mathcal{B}(G)$. Consider the $(\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)})$-graded ring
@@
Multiplication is induced by the tensor product in $\mathscr{K}(G)$. Note that only non-positive shifts $s\leq 0$ produce non-zero homomorphisms. We call $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ the (permutation) twisted cohomology ring of $G$, as in [BG25]. It is convenient to write
@@
  1. (a)
    When $G$ is a finite elementary abelian $p$-group, we recover the twisted cohomology ring constructed in [BG25], since in this case the set of endotrivials $\{u_{N}\}_{N\in\mathcal{N}}$ as defined in [BG25, Example 12.1] and the shift of the tensor unit $k[1]$ form precisely the canonical $\mathbb{Z}$-basis of $\operatorname{Pic}(\mathscr{K}(G))$, as deduced in [MIL24, Section 6].
  2. (b)
    This construction may not immediately adapt for non-$p$-groups, since in this setting it is not known if one has a canonical $\mathbb{Z}$-basis for $\operatorname{Pic}(\mathscr{K}(G))$. A possible replacement we propose could be the subgroup of $\mathscr{K}(G)$ of endotrivials arising from representation spheres for $G$. It is unknown if the subgroup of $\operatorname{Pic}(\mathscr{K}(G))$ spanned by such endotrivials even has maximal rank; we expect these questions to be nontrivial.
To examine the $k$-vector space $\operatorname{H}^{s,q}(G)$, it suffices to choose an indecomposable representative of $k(q)$. This proposition also classifies all the global sections of an effective endotrivial, in the sense of [GAL25].
@@
Let $G$ be a finite $p$-group, $k(q)=C$ an effective endotrivial, and $s$ an integer. Let $f\colon k\to C[s]$ be a nonzero chain complex homomorphism. We have that $f\in\operatorname{Hom}_{\mathscr{K}(G)}(k,C[s])=0$ if and only if $\operatorname{im}(f_{0})\subset C_{s}$ is contained in a contractible direct summand of $C$. In particular, if $C$ is an indecomposable effective endotrivial, then
@@
The converse implication is straightforward. Suppose $f\in\operatorname{Hom}_{\mathscr{K}(G)}(k,C[s])=0$, then there exists a homotopy $h\colon k\to C_{s+1}$ such that $f=d_{s+1}\circ h$. Choose a canonical basis of $C$, and consider the projection of the image of $f$ onto each indecomposable summand. Necessarily on each summand $k[G/H]$ whose projection is nonzero, the image must be the unique minimal nonzero submodule $k\subseteq k[G/H]$. Since $f=d_{s+1}\circ h$, there exists an indecomposable direct summand $M$ of $C_{s+1}$ such that $p_{k[G/H]}\circ f=d_{s+1}\circ p_{M}\circ h$. Since the stabilizers grow condition holds for $C$, the only possible homomorphisms $M\to k[G/H]$ with $k\subset M$ not in the kernel are isomorphisms, hence $M\cong k[G/H]$ and we may split off the contractible summand. An induction argument demonstrates $\operatorname{im}(f)$ is contained in a contractible direct summand of $C$ as desired. The last statement follows from Corollary 3.15. ∎
As in [BAL10a, BG25], we have a canonical comparison map, the heart of the matter. Its existence follows identically to [BG25, Proposition 13.4]. See also [GAL25, Appendix A] for a much more general treatment of these twisted cohomology rings.
@@
Remark 5.5.
Recall for a morphism $f$ in $\mathscr{K}(G)$ we write
The open $\operatorname{open}(f)$ is the preimage of $\operatorname{comp}_{G}$ of the principal quasi-compact open $Z(f)^{c}=\{\mathfrak{p}\mid f\not\in\mathfrak{p}\}$, and is equivalently the open locus of $\operatorname{Spc}(\mathscr{K}(G))$ where $f$ is invertible.
5.1. Twisted cohomology under localization and tt-functors
Next, we’ll extend the results of [BG25, Section 14] regarding localization and functoriality. The following constructions rely on the forerunners constructed in Theorem 4.3.
Definition 5.6.
Let $H$ be a subgroup of $G$. Let $S_{H}\subset\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ be the multiplicative subset generated by all forerunners $\iota^{H}_{C}$, where $C$ runs over the irreducible endotrivials $C\in\mathcal{B}(G)$. We define a $\mathbb{Z}$-graded ring
as the twist-zero part of the localization of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ with respect to $S_{H}$. Explicitly, the homogeneous elements of $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$ consist of fractions $\frac{f}{g}$ where $f,g\in\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ with the same $\mathcal{B}(G)$-twist $q$, and $g$ is a product of “$H$-local isomorphisms” $g^{\prime}\colon k\to C[s]$ with $C\in\mathcal{B}(G)$. Thus, $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$ is $\mathbb{Z}$-graded by the shift only. The homological degree of $\frac{f}{g}$ is the difference sts-t between the shifts of $f$ and $g$.
Construction 5.7.
  1. (a)
    We can perform the central localization (c.f. [BAL10a]) of the whole category $\mathscr{K}(G)$
    $\mathcal{L}_{G}(H):=\mathscr{K}(G)[S_{H}^{-1}].$
    In fact, this localization has idempotent completion $\mathscr{K}(G)|_{U(H)}$ by the same argument as in [BG25, Construction 14.12] or [BAL10a, Theorem 3.6].
    Explicitly, the category $\mathcal{L}_{G}(H)$ is the Verdier quotient of $\mathscr{K}(G)$ by the tt-ideal $\langle\{\operatorname{cone}(g)\mid g\in S_{H}\}\rangle$. It has the same objects as $\mathscr{K}(G)$ and morphisms $x\to y$ of the form $\frac{f}{g}$ where $g\colon k\to C$ belongs to $S_{H}$ for $C$ an effective endotrivial and $f\colon x\to C\otimes y$ is any morphism in $\mathscr{K}(G)$ with the same twist $C$ as the denominator $g$. Moreover, the $\mathbb{Z}$-graded endomorphism ring $\operatorname{End}^{\scriptscriptstyle\bullet}_{\mathcal{L}_{G}(H)}(k)$ of the unit in $\mathcal{L}_{G}(H)$ is the $\mathbb{Z}$-graded ring $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H)$. In particular, $\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(1)\cong\operatorname{H}^{\scriptscriptstyle\bullet}(G)$ by Theorem 4.19. We refer the reader to [BAL10a, Section 3] for this construction in abstract.
  2. (b)
    Twisted cohomology $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is graded over the monoid $\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)}$. Ring homomorphisms induced by tt-functors or localization will be homogeneous with respect to a certain homomorphism $\gamma$ on the corresponding grading monoids.
    Let $H\leq G$ be a subgroup and consider the central localization
    $(-)_{U(H)}\colon\mathscr{K}(G)\twoheadrightarrow\mathcal{L}_{G}(H).$
    Here, the $H$-forerunners all become isomorphisms, yielding a homomorphism on the grading
    $\gamma_{U(H)}\colon\mathbb{Z}\times\mathbb{N}^{\mathcal{B}(G)}\to\mathbb{Z}$
    defined by $\gamma_{U(H)}(s,q)=s+h_{k(q)}(H)$ and one obtains a ring homomorphism
    homogeneous with respect to the monoid homomorphism $\gamma_{U(H)}$.
  3. (c)
    Let $H\trianglelefteq G$ be a normal subgroup. The tt-functor $\Psi^{H}\colon\mathscr{K}(G)\to\mathscr{K}(G/H)$ maps effective endotrivials with h-mark at $G$ equal to 0 to effective endotrivials at $G/H$ equal to 0. In fact, since deflation preserves irreducibility of a representation, $\Psi^{H}$ induces a surjective map $\mathcal{B}(G)\to\mathcal{B}(G/H)\sqcup\{k[1]\}$.
    This defines a homomorphism of graded monoids
    given by $\gamma(s,q)=(s,\overline{q})$ where $\overline{q}$ is given by the surjection $\mathbb{N}^{\mathcal{B}(G)}\twoheadrightarrow\mathbb{N}^{\mathcal{B}(G/H)}$ along the inclusion $\mathcal{B}(G/H)\hookrightarrow\mathcal{B}(G)$ induced from inflation. Therefore, modular fixed points defines a ring homomorphism $\Psi^{H}\colon\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G/H)$ homogeneous with respect to $\gamma_{\Psi^{H}}$. Because $\operatorname{id}_{\mathscr{K}(G/H)}=\Psi^{H}\circ\operatorname{Inf}^{G}_{G/H}$ general algebraic geometry implies $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G/H))$ is homeomorphically embedded into $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ via the map induced by $\Psi^{H}$ on spectra.
  4. (d)
    Given a group homomorphism $\alpha\colon G^{\prime}\to G$, restriction along $\alpha$ also defines a tt-functor $\alpha^{*}:\mathscr{K}(G)\to\mathscr{K}(G^{\prime})$. This again defines a corresponding ring homomorphism $\alpha^{*}\colon\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G^{\prime})$ homogeneous with respect to $\gamma_{\alpha^{*}}$. If $\alpha$ is surjective, then $\alpha^{*}$ is simply inflation $\operatorname{Inf}^{G}_{G/N}$ for $N=\ker(\alpha)$. If $\alpha$ is injective, then $\alpha^{*}$ is restriction. We write
    $\rho^{G}_{H}:=\left(\operatorname{Res}^{G}_{H}\right)^{*}\colon\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)).$
    These are essentially the only two cases one need consider, as every group homomorphism factors as a surjection composed with an injection.
Remark 5.8.
For any subgroup $H\leq G$, the tt-functor $\operatorname{Res}^{G}_{H}\colon\mathscr{K}(G)\to\mathscr{K}(H)$ induces a closed map on Balmer spectra $\operatorname{Spc}(\mathscr{K}(H))\to\operatorname{Spc}(\mathscr{K}(G))$ by [BG25, Proposition 4.7]; this follows by monadicity of the restriction/induction adjunction, see [BAL16]. The same does not hold for the induced map on twisted cohomology $\rho^{G}_{H}\colon\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(H))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$. Indeed, let $G=C_{2}$, then $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(C_{2})\cong k[a,b]$ with $a,b$ the usual $a_{1},b_{1}$ morphisms in $\mathscr{K}(C_{2})$:
[…]
Now consider the restriction $\rho^{C_{2}}_{1}\colon\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(C_{2})\to\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(1)$. The image of $\rho^{C_{2}}_{1}$ is the prime $\langle b\rangle\subset\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(C_{2})$, which is neither an open nor closed point of $\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(C_{2}))$.
Remark 5.9.
Similarly, the localization functor $(-)|_{U(H)}$ obtains the following commutative square.
$U(H)=\operatorname{Spc}(\mathcal{L}_{G}(H))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\mathcal{O}^{\scriptscriptstyle\bullet}_{G}(H))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{\operatorname{comp}_{\mathcal{L}_{G}(H)}}$$\scriptstyle{\operatorname{comp}_{G}}$
Here, the left hand vertical map is the usual graded comparison map from [BAL10a] for the tt-category $\mathcal{L}_{G}(H)$.
We have an analogous version of [BG25, Proposition 14.21].
Proposition 5.10.
Let $H\leq G$ be a subgroup. Then we have a commutative square
If $H\trianglelefteq G$ is normal, then the diagonal is injective.
@@ 5.1. Twisted cohomology under localization and tt-functors
The above diagram factors as follows:
$\operatorname{Spc}(\operatorname{D}_{b}(k[{G}/\!\!/{H}]))$$\operatorname{Spc}(\mathscr{K}({G}/\!\!/{H}))$$\operatorname{Spc}(\mathscr{K}(G))$$\operatorname{Spec}^{h}(\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H}))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}({G}/\!\!/{H}))$$\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$$\scriptstyle{\operatorname{comp}_{\operatorname{D}_{b}(k[{G}/\!\!/{H}])}}$$\scriptstyle{\psi^{H}}$$\scriptstyle{\operatorname{comp}_{{G}/\!\!/{H}}}$$\scriptstyle{\operatorname{comp}_{G}}$$\scriptstyle{\psi^{H}}$
Here, the left-hand square arises from the central localization $(-)|_{U(H/H)}$ which by Theorem 4.19 is identified with the localization $\mathscr{K}({G}/\!\!/{H})\to\operatorname{D}_{b}(k[{G}/\!\!/{H}])$, and the right-hand square is induced from the tt-functor $\Psi^{H}\colon\mathscr{K}(G)\to\mathscr{K}({G}/\!\!/{H})$. Remark 5.9 asserts both diagrams commute, thus the square commutes as well.
If $H\trianglelefteq G$, then $\Psi^{H}\colon\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G/H)$ is surjective, as noted in 5.7(c), hence $\psi^{H}:\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G/H))\hookrightarrow\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$ is injective. Additionally, the comparison map $\operatorname{comp}_{\operatorname{D}_{b}(k[{G}/\!\!/{H}])}$ is always a homeomorphism, thus the bottom left composition is injective, and therefore so is the diagram.
5.2. An identification for the image of the comparison map
We may consider the commutative diagram of 5.10 for all conjugacy classes of subgroups of $G$ at once. Recall the conservative functor of Theorem 1.8, the collection of modular fixed points functors
$\{\hat{\Psi}^{H}\}_{H\in s_{p}(G)/G}\colon\mathscr{K}(G)\to\prod_{H\in s_{p}(G)/G}\operatorname{D}_{b}(k[{G}/\!\!/{H}]).$
For shorthand, we denote this functor by $\hat{\Psi}$. Denote the cohomology ring of $\prod_{H\in s_{p}(G)/G}\operatorname{D}_{b}(k[{G}/\!\!/{H}])$ by
$\operatorname{dH}^{{\scriptscriptstyle\bullet}}(G)=\operatorname{dH}^{{\scriptscriptstyle\bullet}}(G,k)=\prod_{H\in s_{p}(G)/G}\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H},k).$
We have an induced map on cohomology from 5.7, also denoted by $\hat{\Psi}$.
$\hat{\Psi}\colon\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\xrightarrow{\prod\Psi^{H}}\prod_{H\in s_{p}(G)/G}\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}({G}/\!\!/{H})\to\prod_{H\in s_{p}(G)/G}\mathcal{O}_{{G}/\!\!/{H}}^{\scriptscriptstyle\bullet}(H/H)=\operatorname{dH}^{\scriptscriptstyle\bullet}(G).$
We denote the associated map on homogeneous ring spectra by $\hat{\psi}^{h}\colon\operatorname{Spec}^{h}(\operatorname{dH}^{\scriptscriptstyle\bullet}(G))\to\operatorname{Spec}^{h}(\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G))$. The following observation is easy but crucial, as it allows us to understand the comparison map $\operatorname{comp}_{G}$ purely in terms of induced maps on (Zariski) spectra. Injectivity of the diagonal is the crux of Section 6.